A cooper rod of 3mm diameter when subjected to a pull of 495N extends by 0.07mm over a gauge length of 100mm. Calculate Young's modulus for copper


1
Expert's answer
2021-04-14T07:46:42-0400

Young’s modulus = Stress/Strain

Stress = Load /Area of cross section

Area of cross section = πd2/4 = π(3 mm)2/4 = 7.065 mm2

Stress = 495 N / 7.065 mm2 = 70.06 N/mm2

Strain = Stretch / gauge length = 0.07 mm / 100 mm = 0.0007

Young’s modulus = Stress/Strain = 70.06 N/mm2 / 0.0007 = 100 kN/mm2